# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ Appendix D Proofs of Results in Section 5
where we used Lemma A.1 and Assumption 4. By setting $g=f^{\lambda}-f_{0}\in H^{1}(\nu_{X})$ and using Cauchy-Schwarz inequality, we derive
@@
$\leq\lambda\kappa^{1/2}\|\Delta f_{0}+\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}\|f^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}$ (27)
@@
D.1 Deep Sobolev Regressor
This proof use the tenchique of offset Rademacher complexity, proposed by Liang, Rakhlin and Sridharan (2015).
Proof of Lemma 5.2.
@@ D.1 Deep Sobolev Regressor
$(u,v)_{L^{2}({\mathcal{D}})}=\frac{1}{n}\sum_{i=1}^{n}u(X_{i})v(X_{i}),\quad\|u\|_{L^{2}({\mathcal{D}})}^{2}=\frac{1}{n}\sum_{i=1}^{n}u^{2}(X_{i}),\quad\text{for each}~{}u,v\in L^{\infty}(\mu_{X}).$
Then we define the excess risk and its empirical counterpart, respectively, as
@@
    ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\leq c\inf_{\delta>0}\Big{\{}B_{0}^{2}\frac{\operatorname{VCdim}({\mathcal{F}})}{n\delta}+B_{0}\delta\Big{\}},$ (29)
    where $c$ is an absolute positive constant.
@@
    ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+c\inf_{\delta>0}\Big{\{}\sigma^{2}\frac{\operatorname{VCdim}({\mathcal{F}})}{n\delta}+\sigma\delta\Big{\}},$ (30)
    where $c$ is an absolute positive constant.
@@
    $\leq\frac{1}{8}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+2{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}+c\Big{\{}\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)\Big{\}},$
@@
    $\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)=\inf_{f\in{\mathcal{F}}}\Big{\{}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\lambda\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{\}}.$
@@
    ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}\lesssim\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n),$ (32)
@@
    ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}\lesssim\beta\lambda+\frac{\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)}{\lambda}+\frac{\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)}{\lambda}.$ (33)
@@
    $\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)=\inf_{\delta>0}\Big{\{}(B_{0}^{2}+\sigma^{2})\frac{\operatorname{VCdim}({\mathcal{F}})}{n\delta}+(B_{0}+\sigma)\delta\Big{\}}.$
@@
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}R(f)-2\widehat{R}_{{\mathcal{D}}}(f)\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}\Big{]}-\frac{2}{n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}-\frac{2}{n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{3}{2n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}-\frac{1}{2n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}$
$\quad-\frac{3}{2n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}-\frac{1}{2n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{3}{2n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}-\frac{1}{8B_{0}^{2}n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{4}$
$\quad-\frac{3}{2n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}-\frac{1}{8B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}-\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]},$ (34)
@@
where the second inequality follows from the convexity of supremum and Jensen’s inequality, and the third inequality is owing to the fact that $0\leq(f(X_{i})-f_{0}(X_{i}))^{2}\leq 4B_{0}^{2}$ for each $f\in\operatorname{conv}({\mathcal{F}})$.
@@
Let $\delta>0$ and let $\operatorname{conv}({\mathcal{F}})_{\delta}$ be an $L^{2}({\mathcal{D}})$ $\delta$-cover of $\operatorname{conv}({\mathcal{F}})$ satisfying $|\operatorname{conv}({\mathcal{F}})_{\delta}|=N(\delta,\operatorname{conv}({\mathcal{F}}),L^{2}({\mathcal{D}}))$. Then it follows from Cauchy-Schwarz inequality that for each $f\in\operatorname{conv}({\mathcal{F}})$, there exists $f_{\delta}\in\operatorname{conv}({\mathcal{F}})_{\delta}$ such that
$\leq\Big{(}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\Big{)}^{1/2}\Big{(}\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})+f_{\delta}(X_{i})-2f_{0}(X_{i}))^{2}(f(X_{i})-f_{\delta}(X_{i}))^{2}\Big{)}^{1/2}\leq 4B_{0}\delta.$
By a same argument, we obtain
$-\frac{1}{B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}+\frac{1}{B_{0}^{2}n}\sum_{i=1}^{n}(f_{\delta}(X_{i})-f_{0}(X_{i}))^{4}\leq 32B_{0}\delta.$
Combining (34) with above two inequalities yields
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}-20B_{0}\delta$ (35)
$\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\max_{f\in\operatorname{conv}({\mathcal{F}})_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}-\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]}.$
$\operatorname{Pr}_{\varepsilon}\Big{\{}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}>t+\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{\}}.$
For a fixed sample ${\mathcal{D}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, the random variables $\{\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}\}_{i=1}^{n}$ are independent and satisfy ${\mathbb{E}}_{\varepsilon}\big{[}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}\big{]}=0$ and
$-(f(X_{i})-f_{0}(X_{i}))^{2}\leq\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}\leq(f(X_{i})-f_{0}(X_{i}))^{2},$
for each $1\leq i\leq n$. Consequently, it follows from Hoeffding’s inequality (Mohri, Rostamizadeh and Talwalkar, 2018, Lemma D.2) that
$\operatorname{Pr}_{\varepsilon}\Big{\{}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}>t+\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{\}}$
$\leq\exp\Bigg{(}-\frac{(\frac{nt}{3}+\frac{1}{12B_{0}^{2}}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4})^{2}}{2\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}}\Bigg{)}\leq\exp\Big{(}-\frac{nt}{18B_{0}^{2}}\Big{)},$
${\mathbb{E}}_{\varepsilon}\Big{[}\max_{f\in\operatorname{conv}({\mathcal{F}})_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}-\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]}$
[…]
$=T+\frac{18B_{0}^{2}}{n}N(\delta,\operatorname{conv}({\mathcal{F}}),L^{2}({\mathcal{D}}))\exp\Big{(}-\frac{nT}{18B_{0}^{2}}\Big{)}.$
By setting $T=18B_{0}^{2}\log N(\delta,\operatorname{conv}({\mathcal{F}}),L^{2}({\mathcal{D}}))n^{-1}$, we deduces
${\mathbb{E}}_{\varepsilon}\Big{[}\max_{f\in\operatorname{conv}({\mathcal{F}})_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}-\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]}$ (36)
$\leq\frac{18B_{0}^{2}}{n}(1+\log N(\delta,\operatorname{conv}({\mathcal{F}}),L^{2}({\mathcal{D}}))).$
Combining (35) and (36) implies that
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\leq\frac{18B_{0}^{2}}{n}(1+\log N(\delta,\operatorname{conv}({\mathcal{F}}),L^{2}({\mathcal{D}})))+20B_{0}\delta.$ (37)
Using (37) and (A.5) implies (29) directly.
Step (II). Recall the $L^{2}({\mathcal{D}})$ $\delta$-cover $\operatorname{conv}({\mathcal{F}})_{\delta}$ of $\operatorname{conv}({\mathcal{F}})$. There exists $f_{\delta}\in\operatorname{conv}({\mathcal{F}})_{\delta}$ such that $\frac{1}{n}\sum_{i=1}^{n}|f_{\delta}(X_{i})-\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})|^{2}\leq\delta^{2}$, which implies
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f_{\delta}(X_{i}))\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}^{1/2}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}\Big{]}{\mathbb{E}}_{{\mathcal{D}}}^{1/2}\Big{[}\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f_{\delta}(X_{i}))^{2}\Big{]}\leq\sigma\delta,$ (38)
and
$\widehat{R}_{{\mathcal{D}}}^{1/2}(f_{\delta})\leq\Big{(}\frac{1}{n}\sum_{i=1}^{n}(f_{\delta}(X_{i})-\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i}))^{2}\Big{)}^{1/2}+\widehat{R}_{{\mathcal{D}}}^{1/2}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\leq\delta+\widehat{R}_{{\mathcal{D}}}^{1/2}(\widehat{f}_{{\mathcal{D}}}^{\lambda}),$ (39)
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}$ $={\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f_{0}(X_{i}))\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(f_{\delta}(X_{i})-f_{0}(X_{i}))\Big{]}+\sigma\delta$
$\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{\widehat{R}_{{\mathcal{D}}}^{1/2}(\widehat{f}_{{\mathcal{D}}}^{\lambda})+\delta}{\sqrt{n}}\psi(f_{\delta})\Big{]}+\sigma\delta$
$\leq\Big{(}{\mathbb{E}}_{{\mathcal{D}}}^{1/2}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+\delta\Big{)}\frac{1}{\sqrt{n}}{\mathbb{E}}_{{\mathcal{D}}}^{1/2}\Big{[}\psi^{2}(f_{\delta})\Big{]}+\sigma\delta$
$\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+\frac{2}{n}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\psi^{2}(f_{\delta})\Big{]}+\frac{1}{4}\delta^{2}+\sigma\delta.$ (40)
${\mathbb{E}}_{\xi}\Big{[}\psi^{2}(f_{\delta})\Big{]}\leq{\mathbb{E}}_{\xi}\Big{[}\max_{f\in\operatorname{conv}({\mathcal{F}})_{\delta}}\psi^{2}(f)\Big{]}\leq 4\sigma^{2}(\log|\operatorname{conv}({\mathcal{F}})_{\delta}|+1).$ (41)
Combining (40), (41) and Lemma A.5 yields (30).
@@
$=\lambda(\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}+\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-Y_{i})(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f(X_{i}))\leq 0.$
@@
$\lambda(\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}+(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}\leq\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f(X_{i})).$ (42)

### v2

@@ Appendix D Proofs of Results in Section V
where we used Lemma A.1 and Assumption 4. By setting $h=f^{\lambda}-f_{0}\in H^{1}(\nu_{X})$ and using Cauchy-Schwarz inequality, we derive
@@
$\leq\lambda\kappa^{1/2}\|\Delta f_{0}+\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}$
×$\quad\times\|f^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}$ (37)
@@
D-A Deep Sobolev Regressor
Proof of Lemma V.2.
@@ D-A Deep Sobolev Regressor
$(u,v)_{L^{2}({\mathcal{D}})}$ $=\frac{1}{n}\sum_{i=1}^{n}u(X_{i})v(X_{i}),$
$\|u\|_{L^{2}({\mathcal{D}})}^{2}$ $=\frac{1}{n}\sum_{i=1}^{n}u^{2}(X_{i}),$
for each $u,v\in L^{\infty}(\mu_{X})$. Then we define the excess risk and its empirical counterpart, respectively, as
@@
    ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$ (39)
    $\leq 4B_{0}^{2}\Big{\{}\Big{(}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}\Big{)}^{\frac{1}{2}}+\delta\Big{\}}.$
@@
    ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}$ (40)
    $\leq\frac{B_{0}^{2}+\sigma^{2}}{\log^{-1}n}\Big{\{}\Big{(}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}\Big{)}^{\frac{1}{2}}+\delta\Big{\}}.$
@@
    $\leq\frac{1}{8}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+2{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}$
    $\quad+c\Big{\{}\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)\Big{\}},$
@@
    $\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)$
    $=\inf_{f\in{\mathcal{F}}}\Big{\{}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\lambda\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{\}}.$
@@
    ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}$ (42)
    $\lesssim\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n),$
@@
    ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}$ (43)
    $\lesssim\beta\lambda+\frac{\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)}{\lambda}+\frac{\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)}{\lambda}.$
@@
    $\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$
    $=\frac{B_{0}^{2}+\sigma^{2}}{\log^{-1}n}\inf_{\delta>0}\Big{\{}\Big{(}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}\Big{)}^{\frac{1}{2}}+\delta\Big{\}}.$
@@
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}R(f)-\widehat{R}_{{\mathcal{D}}}(f)\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}\Big{]}$
$\quad-\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}$
$\quad-\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\big{(}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}$
$\quad-(f(X_{i})-f_{0}(X_{i}))^{2}\big{)}\Big{]}$
$=2{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$\leq 4B_{0}{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))\Big{]}$
$=4B_{0}{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\Big{]}$
$=4B_{0}{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\Big{]},$ (44)
@@
where the second inequality follows from the convexity of supremum and Jensen’s inequality, and the third inequality holds from Ledoux-Talagrand contraction inequality [97, Lemma 5.7] and the fact that $0\leq|f(X_{i})-f_{0}(X_{i})|\leq 2B_{0}$ for each $f\in\operatorname{conv}({\mathcal{F}})$ and each $1\leq i\leq n$. The last equality invokes the fact that the Rademacher complexity of the convex hull of ${\mathcal{F}}$ is equal to that of ${\mathcal{F}}$.
@@
Let $\delta>0$ and let ${\mathcal{F}}_{\delta}$ be an $L^{2}({\mathcal{D}})$ $(B_{0}\delta)$-cover of ${\mathcal{F}}$ satisfying $|{\mathcal{F}}_{\delta}|=N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))$. Then it follows from Cauchy-Schwarz inequality that for each $f\in{\mathcal{F}}$, there exists $f_{\delta}\in{\mathcal{F}}_{\delta}$ such that
$\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})-\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f_{\delta}(X_{i})$
$\leq\Big{(}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\Big{)}^{1/2}\Big{(}\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-f_{\delta}(X_{i}))^{2}\Big{)}^{1/2}\leq B_{0}\delta.$
Combining (44) with the above inequality yields
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$
$\leq 4B_{0}{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in{\mathcal{F}}_{\delta}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\Big{]}+4B_{0}^{2}\delta$
$\leq 4B_{0}^{2}\Big{(}\frac{2\log|{\mathcal{F}}_{\delta}|}{n}\Big{)}^{\frac{1}{2}}+4B_{0}^{2}\delta$
$=4B_{0}^{2}\Big{(}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}\Big{)}^{\frac{1}{2}}+4B_{0}^{2}\delta,$ (45)
where the last inequality holds from Massart’s lemma [97, Theroem 3.7]. This completes the proof of (39).
Step (II). According to [98, Lemma 4], the Gaussian complexity can be bounded by the Rademacher complexity, that is,
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}f(X_{i})\Big{]}$
$\leq\sigma(\log n){\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\Big{]},$ (46)
where $\varepsilon=\{\varepsilon_{i}\}_{i=1}^{n}$ is a set of Rademacher variables and independent of ${\mathcal{D}}$. By the same argument as (44) and (45), we have
${\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in\operatorname{conv}({\mathcal{F}})}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\Big{]}$
$\leq B_{0}\Big{(}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}\Big{)}^{\frac{1}{2}}+\delta.$ (47)
Combining (46) and (47) completes the proof of (40).
@@
$=\lambda(\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-Y_{i})(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f(X_{i}))\leq 0.$
@@
$\lambda(\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f))_{L^{2}(\nu_{X})}+(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}$ (48)
$\leq\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f(X_{i})).$